# Alchemist: Towards the Design of Efficient Online Continual Learning System

### v1

@@ 8. Related Works
Serving systems:  One of the most popular line of LLM systems research focuses on developing efficient serving systems for LLMs, to improve the serving throughput (Kwon et al., 2023; Anonymous, 2024; Yu et al., 2022; Srivatsa et al., 2024; Zhong et al., 2024a; Patel et al., 2023; Lin et al., 2024a; Cao et al., 2025; Kwon et al., 2023), or reduce the inference latency (Liu et al., 2024; Yao et al., 2024; Miao et al., 2024b; Qiao et al., 2024; Dao et al., 2022). Common optimizations include caching and reusing the KV cache of other prompts (Liu et al., 2024; Gao et al., 2024a; Jin et al., 2024), offloading KV cache to CPU or storage devices (Lee et al., 2024a; Liu et al., 2024; Qin et al., 2025), better parallelization strategy that minimizes network communication overhead (Zhong et al., 2024a; Qin et al., 2025; Microsoft, nd), and faster attention CUDA kernels (Dao et al., 2022; Ye et al., 2025; Lefaudeux et al., 2022). Alchemist is orthogonal to this line of work. The above optimizations developed by previous work can be easily adapted to Alchemist to make its inference part faster and more efficient.
Training systems:  Another line of LLM system work aims at designing efficient systems for LLM training. Some works improve parallelism strategies, such as designing different distributed parallelism for different submodules in multi-modal model training (Huang et al., 2024), designing dynamic parallel strategies according to input sequence length (Li et al., 2024), or reducing idle bubbles in pipeline parallelism (Huang et al., 2019). Other works improve fault-tolerance in distributed training by efficient model checkpointing (Lian et al., 2024; Wan et al., 2024), or quickly detecting faulty machines in distributed training (Deng et al., 2024) and fastly recovers from failure (Wang et al., 2023a). Finally, an emerging line of works build faster training frameworks for RLHF, such as employing intra-stage and inter-stage fusion to improve GPU utilization (Zhong et al., 2024b), or designing APIs for decoupling and orchestrating computation and data dependencies in RLHF dataflows.
@@
GPU multiplexing for training and serving:  Efforts on co-locating inference and training in video analytics relate well to Alchemist. Specifically, Ekya and RECL (Bhardwaj et al., 2020; Khani et al., 2023) enables continuous retraining on new video samples, and jointly host it with inference on the same device by letting them share GPU resource with the technology like Nvidia MPS (nvi, 2025). Co-locating inference and training in LLMs is much harder than that for video analytics because these works assume the workload is deterministic. For example, the resolution and frame rate of the incoming stream is known. Hence, Ekya and RECL understands how much resource (e.g., memory and computation cycles) the training side should occupy.
A more related work, flexLLM (Miao et al., 2024a), co-locates LLM inference and LoRA fine-tuning by fusing serving and training kernels, effectively share the GPU’s computation unit spatially. However, it only considers offline fine-tuning on different input data than that to the serving jobs, thus fails to reuse activations from serving, leading to redundant forward recomputation. This work can be complementary to Alchemist as it’s techniques in kernel fusion can further help Alchemist harvesting the idle computation units even when serving jobs are running.
@@ 9. Conclusion
These findings underscore the potential of co-locating training and serving processes to unlock performance gains in online continual learning scenarios. Moving forward, further optimization of gradient recording techniques and dynamic activation management could pave the way for even more efficient LLM systems, ultimately enabling faster and more scalable AI services.

### v2

@@ 8. Related Works
Serving systems:  One of the most popular lines of LLM systems research focuses on developing efficient serving systems for LLMs, to improve the serving throughput (Kwon et al., 2023; Anonymous, 2024; Yu et al., 2022; Srivatsa et al., 2024; Zhong et al., 2024a; Patel et al., 2023; Lin et al., 2024a; Cao et al., 2025; Kwon et al., 2023), or reduce the inference latency (Liu et al., 2024; Yao et al., 2024; Miao et al., 2024b; Qiao et al., 2024; Dao et al., 2022). Common optimizations include caching and reusing the KV cache of other prompts (Liu et al., 2024; Gao et al., 2024a; Jin et al., 2024), offloading KV cache to CPU or storage devices (Lee et al., 2024a; Liu et al., 2024; Qin et al., 2025), better parallelization strategies that minimize network communication overhead (Zhong et al., 2024a; Qin et al., 2025; Microsoft, nd), and faster attention CUDA kernels (Dao et al., 2022; Ye et al., 2025; Lefaudeux et al., 2022). Alchemist is orthogonal to this line of work. The above optimizations developed by previous work can be easily adapted to Alchemist to make its inference part faster and more efficient.
Training systems:  Another line of LLM system work aims at designing efficient systems for LLM training. Some works improve parallelism strategies, such as designing different distributed parallelism for different submodules in multi-modal model training (Huang et al., 2024), designing dynamic parallel strategies according to input sequence length (Li et al., 2024), or reducing idle bubbles in pipeline parallelism (Huang et al., 2019). Other works improve fault tolerance in distributed training by efficient model checkpointing (Lian et al., 2024; Wan et al., 2024), or quickly detecting faulty machines in distributed training (Deng et al., 2024) and fastly recovering from failure (Wang et al., 2023a). Finally, an emerging line of works builds faster training frameworks for RLHF, such as employing intra-stage and inter-stage fusion to improve GPU utilization (Zhong et al., 2024b), or designing APIs for decoupling and orchestrating computation and data dependencies in RLHF dataflows.
@@
GPU multiplexing for training and serving:  Efforts on co-locating inference and training in video analytics relate well to Alchemist. Specifically, Ekya and RECL (Bhardwaj et al., 2020; Khani et al., 2023) enable continuous retraining on new video samples, and jointly host it with inference on the same device by letting them share GPU resources with technologies like Nvidia MPS (nvi, 2025). Co-locating inference and training in LLMs is much harder than that for video analytics because these works assume the workload is deterministic. For example, the resolution and frame rate of the incoming stream are known. Hence, Ekya and RECL understand how much resource (e.g., memory and computation cycles) the training side should occupy.
A more related work, flexLLM (Miao et al., 2024a), co-locates LLM inference and LoRA fine-tuning by fusing serving and training kernels, effectively sharing the GPU’s computation unit spatially. However, it only considers offline fine-tuning on different input data than that to the serving jobs, thus failing to reuse activations from serving, leading to redundant forward recomputation. This work can be complementary to Alchemist as its techniques in kernel fusion can further help Alchemist harvesting the idle computation units even when serving jobs are running.
@@ 9. Conclusion
These findings underscore the potential of reusing serving activations to unlock performance gains in online continual learning scenarios. Moving forward, further optimization of gradient recording techniques and dynamic activation management could pave the way for even more efficient LLM systems, ultimately enabling more scalable AI services.